1a) objects A and B have the same mass of 3.0kg. They melt when 3.0×10^4 J of heat is added to A and when 9.0×10^4 J is added to B. Determine the latent heat of fusion for the substance from which each object is made.

b) Find the heat required to melt object A when its mass is 6.0kg

2) Find the mass of water that vaporizes when 2.10kg of murcury at 205°C is added to 0.110 kg of water at 80.0°C.
1
Expert's answer
2018-08-14T10:30:08-0400
a) The latent heat of fusion for the substance from which object A is made:
L_a=Q_a/m=(3.0∙〖10〗^4)/3=10 kJ/kg.
The latent heat of fusion for the substance from which object A is made:
L_b=Q_b/m=(9.0∙〖10〗^4)/3=30 kJ/kg.
b)
Q=ML=(10000)(6)=60 kJ.

2) Find the mass of water that vaporizes when 2.10kg of mercury at 205°C is added to 0.110 kg of water at 80.0°C.
Solution
c_m m_m (T_m-T)=c_w m_w (T-T_w )+mL
(140)(2.1)(205-100)=(4186)(0.11)(100-80)+m(2260000)
m=0.00958 kg=9.58 g.